# Towards Fully Autonomous Research Powered by LLMs: Case Study on Simulations Zhihan Liu,a Yubo Chai,a Jianfeng Li,a

### v2

@@
ABSTRACT: The advent of Large Language Models (LLMs) has created new opportunities for the automation of scientific research, spanning both experimental processes and computational simulations. This study explores the feasibility of constructing an autonomous simulation agent (ASA) powered by LLM, through sophisticated API integration, to automate the entire research process, from experimental design, remote upload and simulation execution, data analysis, to report compilation. Using a simulation problem of polymer chain conformations as a case study, we assessed the performance of ASAs powered by different LLMs including GPT-4-Turbo. Our findings revealed that ASA-GPT-4o achieved near-flawless execution on designated research missions, underscoring the potential of LLMs to manage complete scientific investigations autonomously. The outlined automation can be iteratively performed up to twenty cycles without human intervention, illustrating the potential of LLMs for large-scale autonomous research endeavors. Additionally, we discussed the intrinsic traits of ASAs in managing extensive tasks, focusing on self-validation mechanisms and the balance between local attention and global oversight. KEYWORDS: autonomous research, simulation, LLM, polymer physics
@@ INTRODUCTION
In recent years, the rapid development of artificial intelligence technology, particularly the emergence of large language models (LLMs), has brought revolutionary changes to scientific research1, 2, 3, 4, 5. Since the launch of ChatGPT by OpenAI, LLMs such as GPT-4 and Anthropic’s Claude have demonstrated outstanding performance in natural language processing, text generation, and translation 6, 7. These AI models not only excel in traditional text processing tasks but also show significant potential in scientific research, gradually transforming conventional research methods and providing scientists with unprecedented tools and methodologies8, 9, 10, 11.
The application of artificial intelligence (AI) in the fields of chemistry and materials engineering is continually expanding. By integrating machine learning and automated experimental technologies, faster experimental design and data analysis can be achieved, leading to the discovery of new materials and the optimization of existing material properties in a shorter timeframe 12, 13, 14. In the domain of physical chemistry knowledge parsing, Meta’s Galactica assists researchers with tasks such as information organization, knowledge inference, and writing, showcasing the immense potential of large language models in scientific and professional domains 13. Meanwhile, DeepMind’s AlphaFold series of algorithms enable the prediction of structures for biological molecules, from proteins and nucleic acids to larger and more complex complexes 15, 16. In the fields of chemical synthesis and drug design, AI aids in complex chemical synthesis planning, chemical property prediction, and the generation of new materials. Integrating LLMs with specialized chemical tools and databases can extend their capabilities beyond general language tasks 17, 18, 19, 20, 21. Additionally, some studies have highlighted the crucial role of prompt engineering in guiding models to produce accurate and useful outputs. Techniques such as Chain of Thought, ReAct, and Tree of Thoughts help structure the model’s reasoning process, reduce errors, and improve the quality of generated results 20, 22, 23, 24.
Recent research shows that continuous human-AI interaction can accelerate various stages of scientific research. However, this human-led, AI-assisted model still limits the speed and depth of scientific research. Some work has explored an AI-dominated research model, such as smart laboratories (SDLs) that integrate AI, laboratory automation, and robotics to conduct autonomous experiments. These systems can automatically adjust and optimize based on intelligent analysis and decision-making support from experimental results 25, 26, 27. The ultimate goal for future AI scientists is to achieve full automation of scientific research processes, from experimental design to data analysis and conclusion generation. This end-to-end automation is set to become a major development in chemical research and a primary objective in chemical engineering, including both experimental and simulation aspects.
This paper explores the AI automation paradigm in polymer theory simulation research. In traditional simulation research, a supervisor typically discusses and plans the research project with students. The students then write and modify the simulation program based on physical models, run simulations under different parameter conditions on remote servers, collect data, analyze results, and write reports. This paper envisions a new research paradigm where, after the research has been planned, the subsequent tasks are solely carried out by AI. Using an autonomous simulation agent (ASA) powered by LLM, after human researchers input a research plan (RP), the ASA will then take on the role of the student to complete the entire research process. Humans only need to review the AI’s results at the end, thus achieving a research mode without human intervention.
To test this research paradigm, we systematically attempted ASAs powered by different LLMs on a moderately complex simulation problem in polymer physics. This involved an AI fully autonomously coding the simulation program, remote uploading and executing simulations, data organization, plotting, and scientific report writing.
Random walk and self-avoiding walk models are of significant physical importance in describing the spatial configuration of polymer chains. The random walk model assumes that each ‘step’ of the chain is independent and random, while the self-avoiding walk model considers the volume exclusion effects between chain segments. Therefore, they exhibit different dynamic behaviors and scaling properties, with different scaling relationships between the mean square end-to-end distance and the number of chain segmentss N. This scaling relationship has been theoretically derived by Flory, De Gennes, and others 28, 29 and validated through simulations and experiment 30, 31. This problem is fundamental in polymer physics and has been extensively studied. However, for AI, it still presents a novel challenge. The difficulty of predicting the scaling exponent makes it neither too easy nor prohibitively difficult, thus serving as an excellent model problem for testing AI capabilities. We prompted the AI to write simulation programs for both random walk and self-avoiding walk models, compute the mean square end-to-end distance for different chain segment numbers N, fit the results to $\left\langle R^{2}\right\rangle\propto N^{\nu}$ , determine the scaling exponent $\nu$, plot the data, and write a complete scientific report.
The ASA leverages an API automation program (AutoProg) to enable multi-round text interactions with LLMs via API (Figure 1). Initially, the AutoProg provides an RP witten by human researchers to an LLM, and the LLM will break the plan down into multiple sub-tasks and write Python code to complete the first sub-task in its response. The AutoProg then extracts the Python code, executes it, and returns the execution results and any errors along with the conversation history back to the LLM. The LLM then decide whether to modify the Python code or to proceed to the following sub-tasks. This process repeats until the LLM determines that the mission is complete, at which point the AutoProg stops. Upon mission completion, humans evaluate its success based on criteria such as correct simulation, plotting, etc. The ASA we developed is a general system (minor adjustments may be required in the AutoProg for different LLM APIs) that requires only a change in the original RP to be applicable to any simulation problem involving Python programming. In addition to the RPs designed for polymer conformation simulations, we have also crafted RPs for solar system planetary orbit simulations and asteroid trajectory simulations (Figure 1C and SI).
@@
Fig.  1: A. Schematic diagram of the AI automation process for theoretical simulation research conducted in this study. The human researcher writes the research plan (RP) for the theoretical simulation and provides it to the autonomous simulation agent (ASA), which is developed by sophisticated API programming (see Method section). The ASA first generates the simulation program, then uploads it to a remote server, performs simulation calculations with different parameter conditions, collects data, and finally writes a report based on the simulation results. B. Diagram of the ASA dialogue. The human researcher only inputs the RP at the beginning, and then the ASA makes decisions based on the dialogue history, achieving a fully autonomous research mode. C. Two typical simulation problems studied in this paper: the chain conformation simulation of a random walk and the gravitational model simulation of asteroid orbit prediction (See SI).
@@
We compared the performance of ASAs powered by various LLMs, including GPT-4o, and explored common characteristics and issues when LLMs tackle long tasks. These findings provide insights for researchers to better improve and utilize LLMs for AI-driven scientific research.
Overall, this paper presents the first fully autonomous AI-conducted simulation research process powered by LLMs. As demonstrated in the demo video, simply submitting a research plan file via the command ’do AI4S_prj1.txt’ to ASA allows it to execute all subsequent tasks without any human intervention; ASA can even automatically upload programs to a remote server and perform a series of simulations—-something that cannot be achieved solely through interactions with a web-based LLM. By adopting this new research paradigm, we aim to explore the potential of AI in highly complex and data-intensive scientific research, breaking traditional human limitations and advancing the automation and intelligence of scientific research.
@@ Results
Fig.  3: Completion Rates of Various ASAs for RP 1-3. A. Table of LLM models used in this study. B. Statistics of the number of criteria met by each ASA over 20 experiments. We established seven criteria for RP 1-3 to measure mission completion rates and counted the number of times each ASA met these criteria across 20 experiments. Some zero results are not displayed. C. ASA Scoring. Relative scores for each ASA in RP 1-3 were calculated using the EWM and TOPSIS methods. D. ASA Generated Research Report Example. Demonstrates a Word report generated by ASA-GPT-4 for RP 1, including four sections, chain conformation diagrams and a scaling relation fit plot (also see SI).
@@
We designed three simulation missions related to polymer chain conformation modeling to demonstrate and test the ability of ASAs powered by different LLMs to autonomously complete the entire research process. Each mission includes multiple sub-tasks, covering the full spectrum from conducting experiments to analyzing data and writing reports. The research plans (RPs) for these missions are shown in Figure 2. Among these RPs, RP 1 is the simplest involving only basic research steps for simulations given below.
@@
We established several evaluation criteria based on the requirements of the RP, such as whether graphs and plots were generated and whether the derived scaling relationships were correct, to assess task completion. We tested nine ASAs powered by different LLMs, including GPT-4o, conducting 20 trials for each RP and recording the achievement of the evaluation criteria across these trials, as shown in Figure 3.
ASA-GPT-4o and ASA-GPT-4-Turbo excelled on RP 1, achieving near-perfect completion in both simulation and report writing tasks (Figure 3B). ASA-Claude-3.5 and ASA-Qwen-2 followed, performing well in simulation tasks but sometimes generating reports in response or as text files, failing to meet RP 1’s requirements. Other models performed well in plotting graphs but lacked accuracy in establishing correct physical models and deriving correct scaling values through simulations.
Although chain conformation simulations are convenient for comparing different LLMs, they are less convincing due to being previously well-studied topics. To make a stronger case, we need more challenging problems. Additionally, to demonstrate that ASA can solve problems across different scientific domains, we designed RPs for gravitational simulation problems corresponding to RP S1 and RP S2 (Figure S2 and Figure 1C). The mission of RP S2 is to evaluate the risk posed by a meteorite (or asteroid) to earth, given its initial position and velocity and write a report. This mission is challenging even for graduates from physics departments. By simply inputting RP S2 into the same ASA used for chain conformation problems, the ASA will automatically acquire the necessary parameters of the solar system, code the simulation program to calculate the asteroid’s trajectory, and ultimately write the report. This result underscores the effectiveness of our API programming strategy (see Method section) and strongly demonstrates that ASA can address problems in entirely different fields.
@@
RP 2 increases complexity with server interactions and RP 3 doubles the mission length and simulation difficulty. Brief introductions of these two RPs are shown below (Figure 2).
@@
RP 3 required simulating a self-avoiding walk based on the random walk simulation in RP 1. Although the process was longer, it was less complex than RP 2, resulting in improved performance across agents. ASA-GPT-4o, ASA-GPT-4-Turbo, ASA-Claude-3.5, and ASA-Gemini-1.5-Pro performed relatively well. Notably, none of the agents provided a correct self-avoiding scaling due to incorrect sampling methods during simulation, which will be discussed in detail in the Discussion section.
@@
Moreover, ASA-GPT-4o and ASA-Claude-3.5 demonstrated unique advantages in generating rich content. They frequently produced detailed reports exceeding 1,000 words, delving into the distinctions between RW and SA scaling, and occasionally included precise literature citations. ASA-Claude-3.5 referenced accurate self-avoiding walk scaling relationships validated through theoretical derivations and experimental verifications, accompanied by brief analyses of simulation result deviations. Additionally, ASA-GPT-4o presented experimental data in enriched formats, such as frequency distribution plots of end-to-end distances and interactive web-based charts.
Automating Agent Coordination through RP Generation
In the preceding section, RPs were provided by humans. In this section, we task a Main AI with automatically breaking down a human-provided RP into sub-tasks and distributing them as AI RPs to several Subordinate AIs.
We provided RP 1 and 2 directly to the Master AI, adjusting the ASA to ensure each sub-task was handled by a distinct Subordinate AI without access to the Main AI’s or other Subordinate AIs’ conversation histories. The Main AI solely presents the AI RPs (Figure S1A) and received reports upon task completion.
@@
We also employed GPT-4-turbo for this experiment. The results (SI-data-2) revealed that for RP 1, the Main AI effectively decomposes the mission into sub-tasks and successfully distributed them to Subordinate AIs, thereby completing RP 1. However, for RP 2, the Main AI encounters repeated failures. The primary issue lies in precise information transmission. RP 2 provides a random walk simulation program and details of a remote server, including hostname and username. Despite multiple attempts, the Main AI struggles to convey both the simulation program and server details accurately to the Subordinate AIs (Figure S1B). We attempted to instruct the Main AI to include all relevant information in the AI RPs, but this did not notably enhance the accuracy of information transmission. Additionally, we observed instances where AI RPs contained irrelevant information, potentially confusing the Subordinate AIs about the mission’s focus.
@@
Automating the Automation: Crafting Multi-tier RPs for Large-Scale Research
@@
Back in the Automating Basic Simulations section, for each ASA and each RP, we undertake 20 rounds of ‘automation’ by executing RP 1 and collecting the results including codes and files generated by the ASA. Can we automate the entire process of this study on reaserch automation or can we even automate the automation itself?
@@
RP 4 is designed for this purpose. It is nested within RP 1, instructing the Primary AI to execute the command “python AutoProg.py -s p1.txt -n i” 20 times. Each execution generates an Agent AI to fulfill RP 1 contained in p1.txt, while the Primary AI collects all generated files and conducts result analysis. The detailed content of RP 4 is depicted in Figure S1C.
@@
Using ASA-GPT-4-Turbo, we conducted tests on RP 4. The results (SI) strikingly demonstrate that the Primary AI successfully executed the 20 instances of RP 1, meticulously organizing all files into their respective folders and providing a brief summary after all experiments. During this process, the AI independently and without interruption wrote 66 programs (including 26 error versions), conducted more than 20 simulations, generated 120 images, and wrote 20 research reports totaling over 5,000 words. Marvelously, all these missions were accomplished without any human intervention. This illustrates the remarkable capability of multi-tier RP structures in handling large-scale missions.

### v3

# Toward Automated Simulation Research Workflow through LLM Prompt Engineering Design Zhihan Liu,a Yubo Chai,a Jianfeng Li,a

@@
ABSTRACT: The advent of Large Language Models (LLMs) has created new opportunities for the automation of scientific research spanning both experimental processes and computational simulations. This study explores the feasibility of constructing an autonomous simulation agent (ASA) powered by LLMs through prompt engineering and automated program design to automate the entire simulation research process according to a human-provided research plan. This process includes experimental design, remote upload and simulation execution, data analysis, and report compilation. Using a well-studied simulation problem of polymer chain conformations as a test case, we assessed the long-task completion and reliability of ASAs powered by different LLMs, including GPT-4o, Claude-3.5, etc. Our findings revealed that ASA-GPT-4o achieved near-flawless execution on designated research missions, underscoring the potential of methods like ASA to achieve automation in simulation research processes to enhance research efficiency. The outlined automation can be iteratively performed for up to 20 cycles without human intervention, illustrating the potential of ASA for long-task workflow automation. Additionally, we discussed the intrinsic traits of ASA in managing extensive tasks, focusing on self-validation mechanisms, and the balance between local attention and global oversight. KEYWORDS: autonomous research, simulation, LLM, polymer physics
@@ INTRODUCTION
In recent years, the application of artificial intelligence (AI) in scientific research has gained significant traction. AI technologies, through data analysis, pattern recognition, and predictive capabilities, offer researchers unprecedented support, enhancing both efficiency and accuracy.
AI is utilized in chemical synthesis, drug design, and materials science, accelerating the achievement of valuable research outcomes. Meta’s Galactica supports researchers in organizing information, inferring knowledge, and writing, highlighting the potential of AIs in scientific contexts.1 DeepMind’s AlphaFold algorithms enable the prediction of structures for biological molecules, from proteins and nucleic acids to larger and more complex complexes.2, 3 Machine Learning methods not only enhance the understanding and predictive capabilities of complex systems but also offer new approaches to tackling problems that are difficult for traditional computational methods to handle.4, 5, 6, 7, 8, 9 Integrating machine learning with automated experimental techniques enables faster experimental design and data analysis, leading to quicker discoveries of new materials and improvements in existing material properties.10, 11
Among various AI technologies, tools based on large language models (LLMs) such as GPT and Claude have demonstrated superior capabilities in generating and understanding complex texts across a range of natural language processing tasks.12, 13 In addition, LLMs can also generate corresponding code based on text descriptions, significantly simplifying programming tasks.14, 15 Research shows that LLMs excel in handling complex logic, code completion, and debugging, making them valuable assistants in software development and data analysis.16, 17, 18, 19, 20 In the field of scientific research, the application prospects of LLMs are increasingly evident, gradually transforming traditional research methods and providing scientists with unprecedented tools and methodologies.21, 22, 23, 24 For example, integrating LLMs with specialized chemical tools and databases can extend their capabilities beyond general language tasks.25, 26, 27, 28, 29 Researchers have also utilized LLMs to automate literature reviews30, 31, 32, 33 and data analysis34, 35, 36 and assist in experimental design and species screening 37, 38, 39, 29, 40 to improve research efficiency. Additionally, some studies highlighted the crucial role of prompt engineering in guiding models to produce accurate and useful outputs. Techniques such as Chain of Thought, ReAct, and Tree of Thoughts help structure the model’s reasoning process, reduce errors, and improve the quality of generated results.28, 41, 42, 43, 44, 45, 46 These advancements allow researchers to dedicate more time to innovative thinking and theoretical exploration rather than repetitive and inefficient experimental operations.
The aforementioned studies show that continuous human-AI interaction can accelerate various stages of scientific research. However, the current model of these studies mainly involves interaction between humans and AI. By utilizing the robust capabilities of large language models (LLMs), we aspire to gradually reduce human involvement in the research process, aiming to enable AI to independently complete end-to-end workflows, thus further enhancing research efficiency.
Taking simulation research as an example, in traditional simulation research processes, researchers develop a research plan (RP) that specifies the simulation methods to be used, how to write programs or use simulation software, how to build models, and which parameters to test. Following the establishment of the RP, a significant amount of experimentation is carried out to gather data, which are then analyzed to draw scientific conclusions. After the RP is determined, the subsequent steps are largely standardized and present opportunities for automation. Therefore, we intend to test whether, by providing a detailed textual RP, the advanced capabilities of LLMs can allow AI to successfully and automatically complete an entire simulation research process without human intervention. This approach could streamline the research workflow and lead to more efficient outcomes.
With their outstanding language understanding and code generation capabilities, LLMs can produce rich and varied content based on prompts. However, they cannot guarantee the accuracy of the generated code in a single attempt, and due to the length and complexity of tasks described in the RP, LLMs may overlook subtasks. Additionally, as language models, LLMs do not have the ability to control local computers to execute programs. To address these challenges and achieve an automated workflow, we developed an automated system that can identify and run the code generated by the LLM on a local machine. We also used prompt engineering to guide the LLM in producing structured and accurate responses. Through multiple rounds of interaction, we were able to instruct the LLM to debug program errors, check the completeness of task execution, and make the necessary adjustments. This approach resulted in the creation of an end-to-end Autonomous Simulation Agent (ASA). As demonstrated in the demo video, simply submitting an RP txt file via the command “do AI4PS_PRJ1.txt” to ASA allows it to execute all subsequent tasks without any human intervention; ASA can even automatically upload programs to a remote server and perform a series of simulations—something that cannot be achieved solely through interactions with a web-based LLM.
We chose a problem in polymer physics with known outcomes to evaluate the feasibility and reliability of ASA in automating research workflows. Detailed RPs were provided to the ASA, which involved the AI fully autonomously coding simulation programs, remotely uploading and executing simulations, organizing data, plotting results, and writing scientific reports. We also designed several evaluation criteria to determine if the ASA could correctly and completely perform each step according to RP without human intervention.
@@
Fig.  1: (A) Schematic diagram of the AI automation process for theoretical simulation research conducted in this study. The human researcher writes the research plan (RP) for the theoretical simulation and provides it to the autonomous simulation agent (ASA), which is developed by prompt engineering and automated program design (see the Methods section). The ASA first generates the simulation program and then uploads it to a remote server, performs simulation calculations with different parameter conditions, collects data, and finally writes a report based on the simulation results. (B) Diagram of the ASA dialogue. The human researcher only inputs the RP at the beginning, and then, the ASA makes actions based on the given RP and dialogue history, achieving an automated simulation research process. (C) Two simulation problems are used to test the ASA method in this paper: the chain conformation simulation of a random walk and the gravitational model simulation of asteroid orbit prediction (see the SI).
@@
Furthermore, we compared the performance of ASAs powered by various LLMs, including GPT-4o, GPT-4-Turbo, Claude-3.5, and so forth, and explored common characteristics and issues when ASAs tackle long tasks. These findings provide insights for researchers to improve and utilize LLMs better for AI-driven scientific research.
Overall, this study represents the first exploration of an AI-driven, end-to-end simulation research process guided by an RP provided by humans. By using known problems and predefined evaluation criteria, we tested whether methods like ASA, constructed through an LLM prompt engineering and automated program, can accurately and completely execute the entire simulation research process by simply reading the textual information in RP. Through this, we aim to break traditional human limitations and advance the automation and efficiency of scientific research, particularly in data-intensive simulation studies (Figure 1).
@@ Results
Fig.  3: Completion Rates of Various ASAs for RP 1-3. (A) Table of LLM models tested in this study. (B) Statistics of the number of criteria met by each ASA over 20 experiments. We established seven criteria for each RP to measure mission completion rates and counted the number of times each ASA met these criteria across 20 experiments. Some zero results are not displayed. (C) ASA Scoring. Relative scores for each ASA in RP 1-3 were calculated using the EWM and TOPSIS methods (see the Methods section). (D) ASA Generated Research Report Example. Demonstrates screen cuts for a Word report generated by ASA-GPT-4o for RP 1, including four sections, chain conformation diagrams and a scaling relation fit plot ( see the full text in the SI).
Automating Test Simulation Problems
We tested the reliability and stability of the ASA approach using a well-studied polymer chain conformation problem. Random walking and self-avoiding walking models are of significant physical importance in describing the spatial configuration of polymer chains. The random walk model assumes that each “step” of the chain is independent and random, while the self-avoiding walk model considers the volume exclusion effects between chain segments. Therefore, they exhibit different dynamic behaviors and scaling properties, with different scaling relationships between the mean square end-to-end distance and the number of chain segments N. This scaling relationship has been theoretically derived by Flory, De Gennes, and others 47, 48 and validated through simulations and experiments.49, 50 This problem is fundamental to polymer physics and has been extensively studied. However, AI still presents a novel challenge. The difficulty of predicting the scaling exponent makes it neither too easy nor prohibitively difficult, thus serving as an excellent model problem for testing AI capabilities. We hope that the AI can write simulation programs for both random walk and self-avoiding walk models, compute the mean square end-to-end distance $\left\langle R^{2}\right\rangle$ for different chain segment numbers N, fit the results to $\left\langle R^{2}\right\rangle\propto N^{\nu}$, determine the scaling exponent $\nu$, plot the data, and write a complete scientific report.
We designed three simulation missions related to polymer chain conformation modeling to demonstrate and test the ability of ASAs powered by different LLMs to autonomously complete the entire research process based on the given research plans (RPs). Each mission includes multiple subtasks, covering the full spectrum, from conducting experiments to analyzing data and writing reports. The RPs for these missions are listed in Figure 2.
We established several evaluation criteria based on the requirements of RP 1-3, such as whether graphs and plots were generated and whether the derived scaling relationships were correct, to assess task completion. We tested nine ASAs powered by different LLMs, including GPT-4o, conducting 20 trials for RP 1-3 and recording the achievement of the evaluation criteria across these trials, as shown in Figure 3. Subsequent sections will analyze the performance of ASAs on RPs 1-3.
Despite well-studied chain conformation simulations being able to effectively demonstrate the reliability of ASA in automatically executing simulation research procedures based on RPs, to demonstrate that ASA can solve problems across different scientific domains, we also designed RPs for more challenging gravitational simulation problems corresponding to RP S1 and RP S2 (Figure S2 and 1C). The mission of RP S2 is to evaluate the risk posed by a meteorite (or asteroid) to Earth, given its initial position and velocity and write a report. This mission is challenging even for graduates of physics departments. By simply inputting RP S2 into the same ASA used for chain conformation problems, the ASA will automatically acquire the necessary parameters of the solar system, code the simulation program to calculate the asteroid’s trajectory, and ultimately write the report. This result underscores the effectiveness of our prompt engineering and automated program design strategy (see the Methods section) and strongly demonstrates that ASA can address problems in entirely different fields.
@@
Among RP 1-3, RP 1 is the simplest, involving only basic research steps for simulations given below.
@@
ASA-GPT-4o and ASA-GPT-4-Turbo excelled on RP 1, achieving near-perfect completion in both simulation and report writing tasks (Figure 3B). ASA-Claude-3.5 and ASA-Qwen-2 followed, performing well in simulation tasks but sometimes generating reports in response or as text files, failing to meet RP 1’s requirements. Other models performed well in plotting graphs but lacked accuracy in establishing the correct physical models and deriving correct scaling values through simulations.
@@
RP 2 increases complexity with server interactions, and RP 3 doubles the mission length and simulation difficulty. Brief introductions of these two RPs are shown below (Figure 2).
@@
RP 3 required simulating a self-avoiding walk based on the random walk simulation in RP 1. Although the process was longer, it was less complex than RP 2, resulting in an improved performance across agents. ASA-GPT-4o, ASA-GPT-4-Turbo, ASA-Claude-3.5, and ASA-Gemini-1.5-Pro performed relatively well. Notably, none of the agents provided a correct self-avoiding scaling due to incorrect sampling methods during simulation, which will be discussed in detail in the Discussion section.
@@
Moreover, ASA-GPT-4o and ASA-Claude-3.5 demonstrated unique advantages in generating rich content. They frequently produced detailed reports exceeding 1000 words, delving into the distinctions between RW and SA scaling, and occasionally included precise literature citations. ASA-Claude-3.5 referenced accurate self-avoiding walk scaling relationships validated through theoretical derivations and experimental verifications, accompanied by brief analyses of simulation result deviations. Additionally, ASA-GPT-4o presented experimental data in enriched formats, such as frequency distribution plots of end-to-end distances and interactive web-based charts.
Automating Agent Coordination through Sub-RP Generation
In the preceding section, the ASA used a single AI to execute all of the content provided in the RP given by humans. In this section, we adjusted the logic for storing and managing dialogue history with the LLM within the ASA. We introduced a Manager AI that automatically breaks down a human-provided RP into subtasks and distributes them as sub-RPs to multiple Executor AIs. Each Executor AI is unaware of the dialogue history of the Manager AI and other Executor AIs, focusing solely on executing the assigned sub-RP. In contrast, the Manager AI solely presents the sub-RPs (Figure S1A) and receives reports upon task completion.
@@
We used GPT-4-Turbo for the experiments, testing ASA’s completion of RP 1 and RP 2 using the Manager-Executor AI mode. The results (SI-data-2) revealed that for RP 1, the Manager AI effectively decomposes the mission into subtasks and successfully distributes them to Executor AIs, thereby completing RP 1. However, for RP 2, the Manager AI encounters repeated failures. The primary issue lies in precise information transmission. RP 2 provides a random walk simulation program and details of a remote server, including the hostname and username. Despite multiple attempts, the Manager AI struggles to convey both the simulation program and server details accurately to the Executor AIs (Figure S1B). We attempted to instruct the Manager AI to include all relevant information in the sub-RPs, but this did not notably enhance the accuracy of the information transmission. Additionally, we observed instances where sub-RPs contained irrelevant information, potentially confusing the Executor AIs about the mission’s focus.
@@
Automating the Automation: Crafting Multitier RPs for the Large-Scale Research Process
@@
Back in the Automating Basic Simulations section, for each ASA and each RP, we undertake 20 rounds of “automation” by executing RP 1 and collecting the results, including codes and files generated by the ASA. Can we automate the entire process of this 20-round test on research automation, or can we automate the process of automation itself?
@@
RP 4 is designed for this purpose. It is nested within RP 1, instructing a Primary AI to execute the command “python AutoProg.py -s p1.txt -n” 20 times. Each execution generates a Secondary AI to fulfill RP 1 contained in p1.txt, while the Primary AI collects all generated files and conducts result analysis. The detailed content of RP 4 is depicted in Figure S1C.
@@
Using ASA-GPT-4-Turbo, we conducted tests on RP 4. The results (SI) strikingly demonstrate that the Primary AI successfully executed the 20 instances of RP 1, meticulously organizing all files into their respective folders and providing a brief summary after all experiments. During this process, ASA-GPT-4-Turbo independently and without interruption wrote 66 programs (including 26 error versions), conducted more than 20 simulations, generated 120 images, and wrote 20 research reports totaling over 5000 words. Marvelously, all of these missions were accomplished without any human intervention. This illustrates the remarkable capability of multitier RP structures in handling large-scale missions.
Promoting AI’s Autonomy—Brief-RP ASA
To enhance the autonomy and flexibility of the AI, we developed the brief-RP ASA. We provide the system with a concise description of the research plan (RP) in the form of one or two sentences, allowing the AI to generate a detailed RP, including parameter settings, and subsequently execute the plan (see Section II in the SI for further details). Using brief-RP ASA, we addressed three distinct research questions (Figures S4-S6), such as predicting the viscosity of polymers based on a given data set using two machine learning methods and investigating the impact of natural disaster frequency on the dynamics of a simple grassland ecosystem. These examples illustrate that AI powered by large language models (LLMs) can exhibit a degree of creativity and autonomy in simulation-based research and that our ASA can be effectively applied to a wide range of research problems. While ASA was able to produce valuable insights in ecosystem studies, such as identifying a threshold above which increased disaster frequency leads to extinction events, its ability to generate entirely novel knowledge remains limited, and the reliability of its results requires further validation. Nevertheless, as LLM capabilities continue to advance, the potential of brief-RP ASA is boundless, enabling the rapid execution of a large number of simulation tasks.